OpenStudy (anonymous):

The formula r= 2sqrt5L can be used to approximate the speed r, in miles per hour of a car that has left skid marks of length L, in feet. How many feet will a car skid at 70 mph?

jimthompson5910 (jim_thompson5910):

\[r=2\sqrt{5L}\] \[70=2\sqrt{5L}\] \[\frac{70}{2}=\sqrt{5L}\] \[35=\sqrt{5L}\] \[35^2=5L\] \[1225=5L\] \[\frac{1225}{5}=L\] \[245=L\] \[L=245\]

OpenStudy (anonymous):

\[r=2\sqrt{5L}\]

OpenStudy (anonymous):

so would the answer be 1225 feet?

jimthompson5910 (jim_thompson5910):

answer would be 245 ft

OpenStudy (anonymous):

oops